A rowing team consists of four rowers who weigh 152, 156, 160 and 164 pounds. 

 

  1. How many samples of size n=2 can be drawn from this population ?
  2. Construct a table showing all possible samples, mean and probability of each sample.
  3. Construct the sampling distribution ( probability distribution ) of the sample means.
  4. Draw the histogram of the sampling distribution of the means
1
Expert's answer
2022-04-26T17:44:43-0400

1. "C^2_4=\frac{4!}{2!2!}=6"

2. m(152,156)=(152+156)/2=154


m(152,160)=(152+160)/2=156


m(152,164)=(152+164)/2=158


m(156,160)=(156+160)/2=158


m(156,164)=(156+164)/2=160


m(160,164)=(160+164)/2=162

3. Frequencies

F(154)=F(156)=F160)=F(162)=1

F(158)=2

Probabilities "P(x)=F(x)/\sum F(x)"

P(154)=P(156)=P(160)=P(162)=1/6

P(158)=2/6